# From Reliability to Security: How RIS-Assisted Adaptive SM and SSK Enhances Wireless Systems

### v1

@@ VII Conclusion
This paper presented two novel RIS-assisted IM schemes, RASM and RASSK, aimed at enhancing SE and ensuring PLS in wireless transmissions. A comprehensive analysis of SE, BER, detection complexity, and SR confirms that the proposed schemes provided reliable communication quality and protect confidential information from eavesdropping. Compared to traditional IM schemes, they offered better SE, PLS performance, and robustness against imperfect CSI. To further approach practical applications, more sophisticated eavesdropper model will be considered in our future works. Overall, these schemes showed strong potential for future wireless applications.
@@
The CSI estimation at receiver can hardly be perfect especially in complex wireless environment. For simplicity, the imperfect CSI estimation is only considered in the channel from the RIS to receive. According to this and referring to the imperfect CSI model from [31]-[32], the coefficient of $\check{h}_{2}^{i,m}$ can be expressed as the sum of $\dot{h}_{2}^{i,m}\sim\mathcal{C}\mathcal{N}(0,1-\delta_{e1}^{2})$ and $\ddot{h}_{2}^{i,m}\sim\mathcal{C}\mathcal{N}(0,\mathrm{}\delta_{e1}^{2})$. Giving that $\delta_{e2}^{2}$ denotes the inaccuracy of estimation in CSI with $\ddot{\delta_{e2}^{2}}=1-\delta_{e2}^{2}$. Thus, we can respectively obtain the receive signals at $l$-th receive antenna in $r$-th selected AC for the RASM and RASSK schemes with imperfect CSI estimation as
$\ddot{y}_{l}^{RASM}=\underset{\mathrm{constructive}\,\,part}{\underbrace{\sum_{i=\left(l-1\right)N_{E}+1}^{N_{E}}{\left(\delta_{e2}^{2}\dot{h}_{2}^{i,m}+\ddot{\delta}_{e2}^{2}\ddot{h}_{2}^{i,m}\right)\Phi_{i,l}h_{1}^{i}}x_{k}}}+$
@@
$\ddot{y}_{l}^{RASSK}=\underset{\mathrm{constructive}\,\,part}{\sqrt{E_{s}}\underbrace{\sum_{i=\left(l-1\right)N_{E}+1}^{N_{E}}{\left(\delta_{e2}^{2}\dot{h}_{2}^{i,m}+\ddot{\delta}_{e2}^{2}\ddot{h}_{2}^{i,m}\right)\Phi_{i,l}h_{1}^{i}}}}+$

### v2

@@ VII Conclusion
This paper presented two novel RIS-assisted IM schemes, RASM and RASSK, aimed at enhancing SE and ensuring PLS in wireless transmissions. When the number of receive antennas is 8 or more, the proposed scheme achieves approximately twice the SE of the traditional RSM scheme, while only providing an SNR loss of around 0.5 dB at a target BER lower than $10^{-3}$. Moreover, our physical layer security evaluations confirm that the spatial channel decorrelation inherent to our design strictly guarantees a positive secrecy rate, thoroughly validating the system’s robustness against unauthorized eavesdropping. To further approach practical applications, more sophisticated eavesdropper models, as well as the evaluation of the proposed schemes in dynamic multi-user environments with high mobility, will be considered in our future works. Overall, these schemes showed strong potential for future wireless applications.
@@
The CSI estimation at receiver can hardly be perfect especially in complex wireless environment. For simplicity, the imperfect CSI estimation is only considered in the channel from the RIS to receive. According to this and referring to the imperfect CSI model from [40, 41], the coefficient of $\check{h}_{2}^{i,m}$ can be expressed as the sum of $\dot{h}_{2}^{i,m}\sim\mathcal{C}\mathcal{N}(0,1-\delta_{e1}^{2})$ and $\ddot{h}_{2}^{i,m}\sim\mathcal{C}\mathcal{N}(0,\mathrm{}\delta_{e1}^{2})$. Giving that $\delta_{e2}^{2}$ denotes the inaccuracy of estimation in CSI with $\ddot{\delta_{e2}^{2}}=1-\delta_{e2}^{2}$. Thus, we can respectively obtain the receive signals at $l$-th receive antenna in $r$-th selected AC for the RASM and RASSK schemes with imperfect CSI estimation as
$\ddot{y}_{l}^{RASM}=\underset{\mathrm{constructive}\,\,part}{\underbrace{\sum_{i=\left(l-1\right)N_{E}+1}^{zN_{E}}{\left(\delta_{e2}^{2}\dot{h}_{2}^{i,m}+\ddot{\delta}_{e2}^{2}\ddot{h}_{2}^{i,m}\right)\Phi_{i,l}h_{1}^{i}}x_{k}}}+$
@@
$\ddot{y}_{l}^{RASSK}=\underset{\mathrm{constructive}\,\,part}{\sqrt{E_{s}}\underbrace{\sum_{i=\left(l-1\right)N_{E}+1}^{zN_{E}}{\left(\delta_{e2}^{2}\dot{h}_{2}^{i,m}+\ddot{\delta}_{e2}^{2}\ddot{h}_{2}^{i,m}\right)\Phi_{i,l}h_{1}^{i}}}}+$
@@
In the meanwhile, the received signals with imperfect CSI estimation of $u$-th unselected received antenna in the RASM and RASSK schemes can be respectively given as:
$\ddot{y}_{u}^{RASM}$
$=\sum_{q=1}^{N_{a}}{\sum_{i=\left(q-1\right)N_{E}+1}^{qN_{E}}{\left(\delta_{e2}^{2}\dot{h}_{2}^{i,m}+\ddot{\delta}_{e2}^{2}\ddot{h}_{2}^{i,m}\right)\Phi_{i,l}h_{1}^{i}}}x_{k}+\tau_{u},$ (84)
$\ddot{y}_{u}^{RASSK}$
$=E_{s}\sum_{q=1}^{N_{a}}{\sum_{i=\left(q-1\right)N_{E}+1}^{qN_{E}}{\left(\delta_{e2}^{2}\dot{h}_{2}^{i,m}+\ddot{\delta}_{e2}^{2}\ddot{h}_{2}^{i,m}\right)\Phi_{i,l}h_{1}^{i}}}+\tau_{u}.$ (85)